OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (-3, -4) and (0, 2)

OpenStudy (uri):

First find slope

Elsa213 (elsa213):

\(\large{\frac{\Delta y}{\Delta x}= \frac{y_2-y_1}{x_2-x_1}}\)

OpenStudy (anonymous):

Yeah what @uri said

OpenStudy (anonymous):

(y + 3) = -2(x + 4) right?

OpenStudy (anonymous):

I doing the math now

OpenStudy (anonymous):

Slope is -2

OpenStudy (anonymous):

So that is the correct answer

OpenStudy (anonymous):

woohoo :D thanks ! i wasn't sure at first i had the right answer.

OpenStudy (anonymous):

If i was help medal please

OpenStudy (anonymous):

No problem anytime